# DatAasee: A Metadata-Lake as Metadata Catalog for a Virtual Data-Lake

### v1

@@ 5.2.3 Presentation-Tier
As an API-based web-service, cf. Data Commons (Grossman et al. 2016), the primary front-end of the DatAasee metadata-lake is a REST-like (REpresentational State Transfer) API (Application Programming Interface) with CQRS (Command-Query-Responsibility-Segregation) aspects over HTTP. While the ingest encoding is XML, the HTTP-API only accepts and responds exclusively in JSON. As a secondary effect, DatAasee can also serve as translator from legacy XML sources to JSON. The response content is formatted following the JSON-API (Katz et al. 2022) standard. And to establish self-documenting messages, each response contains a link to its content’s associated JSON Schema, such that the wrapping JSON-API and its payload data are understandable. Furthermore, this API is documented via an OpenAPI (Miller et al. 2020) document which is provided by the API itself.
@@ 6 Preliminary Evaluation
Compatibility No Only with DataCite and FRBR
@@ 6.1 FAIR as Alternative Features
Instead of the previous features, an alternative set of features for the evaluation of metadata layers or metadata systems can be derived from the FAIR principles [^8] See also: https://www.go-fair.org/fair-principles. Those are explicitly not meant to be added to the list of existing features, but rather as an alternative, since their coverage of underlying qualities overlaps with the functional, generic, and semantic features above. Yet overall, the following criteria are more concerned with the outward functionality than internal structure:
  • Findability is given by the exposed search modes, such as full-text (F1), facets (F2), or via query languages (F3).
@@ 7 Summary
A foundational attribute for research data, research software, or publications to be useful is first and foremost their availability (Fehr et al. 2016). So, a data-lake - or metadata-lake - with its priority of load before transform (in the context of ELT) qualifies for the task of research metadata management.
The metadata cataloging aspect of the metadata-lake then makes it a suitable metadata layer for a virtual data-lake or FAIR-compliant repository.
DatAasee is an open-source implementation of the proposed metadata-lake architecture, and aims to become the metadata-layer for academic institutions such as universities and research institutes, as well as academic libraries, scientific libraries or research libraries.
@@
The source code for DatAasee (version 0.2) is available under the MIT-license at:
Appendix
Below is the DatAasee data-tier database schema, encoded in yasql, visualized:

### v3

@@ 5.2.3 Presentation-Tier
As an API-based web-service, cf. Data Commons (Grossman et al. 2016), the primary front-end of the DatAasee metadata-lake is a REST-like (REpresentational State Transfer) API (Application Programming Interface) with CQRS (Command-Query-Responsibility-Segregation) aspects over HTTP. While the ingest encoding is XML, the HTTP-API only accepts and responds exclusively in JSON. As a secondary effect, DatAasee can also serve as translator from legacy XML sources to JSON. The response content is formatted following the JSON-API (Katz et al. 2022) standard. And to establish self-documenting messages, each response contains a link to its content’s associated JSON Schema, such that the wrapping JSON-API and its payload data are understandable. Furthermore, this API is documented via an OpenAPI (Miller et al. 2020) document which is provided by the API itself. Since all messages to and from the DatAasee API server are in JSON encoding, for all requests and responses JSON Schemas are provided through the API itself.
@@ 6 Preliminary Evaluation
Compatibility No Only with DataCite and OpenWEMI
@@ 6.1 FAIR as Alternative Features
An alternative set of features for the evaluation of metadata layers or metadata systems can be derived from the FAIR principles [^8] See also: https://www.go-fair.org/fair-principles. Those are not meant to be added, but rather as an alternative, since their coverage of underlying qualities overlaps with the features above. In comparison, the following criteria are concerned with the outward functionality instead of internal structure:
  • Findability is given by the exposed search modes, such as full-text (F1), filters or facets (F2), or via query languages (F3).
@@ 7 Summary
A foundational attribute for research data, research software, or publications to be useful is first and foremost their availability (Fehr et al. 2016). So, a data-lake - or metadata-lake - with its priority of load before transform (in the context of ELT) qualifies for the task of research metadata management. The metadata cataloging aspect of the metadata-lake then makes it a suitable metadata layer for a virtual data-lake, dataspace, or FAIR-compliant repository.
DatAasee is an open-source implementation of the proposed metadata-lake architecture, and aims to become the metadata-layer in academic institutions such as universities and research institutes. Particularly, to automate metadata handling and improve research data discoverability in university libraries, academic libraries, scientific libraries or research libraries.
@@
The source code for DatAasee (version 0.5) is available under MIT-license at:
@@ Acknowledgement
Appendix
Below is the DatAasee data-tier database schema, encoded in yasql [^9] See: github.com/aryelgois/yasql, visualized: